You are conducting a multinomial hypothesis test for the claim that the 4 categories occur with the following frequencies:

HoHo : pA=0.15pA=0.15;  pB=0.3pB=0.3;  pC=0.4pC=0.4;  pD=0.15pD=0.15


Category Observed
Frequency
A 39
B 53
C 55
D 26



What is the chi-square test-statistic for this data?
χ2=χ2= Incorrect
Answer #1

Following table shows the calculations:

Category O E-p*173(O-E)A2/E 39 0.1525.95 6.562716763 51.9 0.023314066 69.2 2.913872832 0.1525.95 9.63391E-05 9.5 0.3 0.4 53 2

The test statistics is

\chi^{2}=\sum \left ( \frac{(O-E)^{2}}{E} \right )=9.5